An iron rod has a length of 70m when the temperature is 30°C. By how much will it expand when the temperature rises to 40°C (Linear expansion of iron rod = 1.2×10^-5 per kelvin)


1
Expert's answer
2022-02-18T15:53:49-0500

Given:

"l_0=70\:\rm cm"

"\alpha=1.2*10^{-5}\:\rm K^{-1}"

"\Delta t=10\:\rm K"


The thermal expansion is described by an equation



"\Delta l=l_0\alpha\Delta t""\Delta l=\rm 70\: cm*1.2*10^{-5}\: K^{-1}*10\: K=0.0084\: cm"